If Jan can paint a house in 4 hours and Bill can paint the same house in 6, how long will it take for both of them to paint the house together...?
Trying to figure out how to set this one up, it should be simple and was thinking of converting into a fraction with a common denominator... but was sure if that was making it harder than it was ? thanks
Answer given is 2 hrs 24 mins just not sure how 2 get there....

Answer by nerdybill(7384)

If Jan can paint a house in 4 hours and Bill can paint the same house in 6, how long will it take for both of them to paint the house together...?
Trying to figure out how to set this one up, it should be simple and was thinking of converting into a fraction with a common denominator... but was sure if that was making it harder than it was ? thanks
Answer given is 2 hrs 24 mins just not sure how 2 get there....
.
Let x = time it takes to paint the house together
then
x(1/4 + 1/6) = 1
Multiplying both sides by a common denominator 24:
24x(1/4 + 1/6) = 24(1)
x(6 + 4) = 24
x(10) = 24
x = 2.4 hours
Or, converting to hours and minutes:
x = 2 hours and (.4*60) mins
x = 2 hours and 24 mins